# Secure Human Oversight of AI: Exploring the Attack Surface of Human Oversight

### v1

@@ Abstract
Human oversight of AI is promoted as a safeguard against risks such as inaccurate outputs, system malfunctions, or violations of fundamental rights, and is mandated in regulation like the European AI Act. Yet debates on human oversight have largely focused on its effectiveness, while overlooking a critical dimension: the security of human oversight. We argue that human oversight creates a new attack surface within the safety, security, and accountability architecture of AI operations. Drawing on cybersecurity perspectives, we analyze attack vectors that threaten the requirements of effective human oversight, thereby undermining the safety of AI operations. Such attacks may target the AI system, its communication with oversight personnel, or the personnel themselves. We then outline hardening strategies to mitigate these risks. Our contributions are: (1) introducing a security perspective on human oversight, and (2) providing an overview of attack vectors and hardening strategies to enable secure human oversight of AI.
@@ 1 Introduction
Figure 1: The human oversight attack surface encompasses the AI system’s technical infrastructure, the human oversight personnel, and the communication channels between the two sides. Attacks on human oversight aim to undermine oversight personnel’s epistemic access, causal power, self-control, and fitting intentions, the four requirements of effective human oversight [1]. Malicious actors can utilize different attack vectors to undermine one or several of these requirements (see Table 1). In order to protect human oversight from malicious actors, various hardening strategies need to be considered and deployed to counteract the threats posed by the attack vectors (see Table 2). A well-designed combination of hardening strategies can shield against different attack vectors and, thus assure secure human oversight.
@@
Human oversight of artificial intelligence (AI) is often considered a strategy to address risks associated with the use of AI in high-risk contexts [2, 3, 4]. Human oversight personnel are expected to detect inaccurate outputs, outputs that may pose risks to fundamental human rights (e.g., discriminatory outputs), unsafe states of AI systems, and system malfunctions, and to intervene in order to mitigate or minimize these risks [5]. Human oversight is also a key concept in AI governance and ethics [3, 2] and a central requirement for the implementation of high-risk AI systems in emerging AI regulation, such as Article 14 of the European AI Act [6, 3]. However, the idea of relying on humans to oversee AI in order to mitigate operational risks also faces criticism and is expected to encounter substantial challenges in practice [7, 2, 1]. The most prominent criticisms and challenges include doubts about whether humans can effectively oversee AI systems [2], the difficulty of defining what constitutes effective human oversight [1], clarifying who should provide oversight and when it should occur [6], specifying which risks and unwanted system behaviors oversight personnel should detect [5], and determining how to design and evaluate effective human oversight [1].
@@
However, one key challenge that remains underexplored is the security risks that arise from human oversight of AI. Because human oversight will be implemented as part of the safety, security, and accountability architecture of AI systems [3, 7], malicious actors who successfully attack elements related to oversight can significantly undermine the safety of AI operations. In this sense, the human oversight architecture creates a new attack surface for targeting AI operations in high-risk contexts such as critical infrastructure, public administration, or medicine. A security-focused perspective on human oversight is crucial, given the central role oversight already plays, the expectations placed on it in high-risk contexts, and its growing importance as a regulatory requirement in AI governance, such as under the EU AI Act. If we fail to assess and address the security challenges of human oversight, we risk overlooking potential attack vectors in high-risk contexts; and without awareness of these vectors, we will not be able to design effective hardening strategies.
In this paper, we introduce a security-related perspective on human oversight of AI. Specifically, drawing on cybersecurity research, we outline possible attack vectors that threaten the key requirements of effective human oversight (i.e., epistemic access, self-control, causal power, and fitting intentions; [1]). These attacks may target the AI system itself, the communication between the AI system and oversight personnel, or the personnel directly (see Figure 1). We also draw on the cybersecurity literature to discuss hardening strategies that can mitigate the risks posed by such attacks.
@@
This paper makes two key contributions. First, we introduce a security-related perspective on human oversight of AI. This perspective is critically missing from current research, even though human oversight is becoming an integral component of the safety, security, and accountability architecture of AI operations, and thus an increasingly viable target for malicious actors. Second, we provide an overview of possible attack vectors and hardening strategies contextualized for human oversight of AI. We hope this contribution supports AI governance and system design in ensuring not only effective but also secure human oversight of AI.
@@ 2.1 Effective Human Oversight of AI
Research discusses the nature of human oversight of AI, including the what, when, and who of oversight. Regarding the what, many forms of human-AI collaboration in principle involve some degree of human oversight. Across a wide range of human-AI interaction use cases, humans monitor AI system outputs, check their accuracy, request explanations, adapt outputs, intervene when systems malfunction, or take over manual control in the event of failures [8, 9, 10, 11, 12, 13, 14, 15]. In a more targeted sense, and following how emerging research has tentatively defined the concept [1], human oversight specifically refers to humans mitigating risks in AI operations [16, 6]. More concretely, oversight personnel are tasked with monitoring AI systems, detecting failures and malfunctions, identifying unsafe states, flagging inaccurate outputs, and recognizing outputs that may pose risks to fundamental human rights, such as discriminatory outputs [1, 5, 6]. Once such issues are detected, oversight personnel are expected to act in ways that effectively mitigate emerging risks [5, 1]. This can involve taking over manual control, overwriting inaccurate or inadequate outputs [17], or stopping AI operations and triggering intervention by other human stakeholders, such as developers who can adapt the system. Importantly, the AI system under oversight may be operated by another human who is not considered oversight personnel but a user of the system. For example, a patient using an AI chatbot for mental health support [18] would be the user, while oversight personnel monitor the patient–AI interaction to detect risky situations (e.g., if the patient discusses suicidal behavior with the chatbot).
@@
Regarding the when of human oversight, human involvement is important at every stage of AI development: during system design, at run time, and when inspecting errors and failures in hindsight [4, 19, 20]. In a more targeted sense, human oversight usually refers to personnel overseeing AI systems during run time [1, 3]. This can involve dynamic and time-critical situations, such as humans overseeing autonomous vehicles [17, 21], but also less dynamic contexts, such as AI in judicial decision-making or hiring, where there is more time to deliberate about the accuracy or adequacy of AI outputs [5, 22].
@@
Regarding the who of human oversight, any human with the skills necessary to effectively oversee an AI system could serve as oversight personnel [6, 23]. These individuals may be domain experts such as doctors or judges, who can evaluate the adequacy of specific AI outputs based on their expertise. Alternatively, oversight personnel may be trained to detect AI failures without being domain experts in the task performed by the AI. For example, in a medical AI setting, one oversight person might monitor the system’s technical functions on an aggregate level, while medical professionals evaluate individual outputs requiring domain knowledge [23, 5]. Human oversight can also be performed by a single individual or by a team collaboratively overseeing AI systems.
@@
Research has also examined the requirements for effective human oversight. Sterz et al. [1] identify four key aspects: epistemic access, causal power, self-control, and fitting intentions. As we discuss possible attack vectors, we consider how successful attacks might undermine these requirements. Epistemic access means that oversight personnel possess sufficient knowledge about the AI system and understand the consequences of their interventions. Without it, they may fail to detect system failures or inaccurate outputs. Causal power refers to the ability to intervene in ways that actually influence AI processes and outputs. Without causal power, oversight personnel may be unable to overwrite or correct system outputs. Self-control requires oversight personnel to maintain agency over their own actions. For instance, fatigue or intoxication may impair their capacity to fulfill oversight duties. Likewise, being threatened by another person can also undermine self-control, as the actions performed by the oversight personnel are then steered by another agent. Fitting intentions means that personnel are motivated to carry out their responsibilities with the right intentions. In an extreme case, a malicious actor might be hired as oversight personnel and deliberately undermine AI safety. In a less extreme case, poor working conditions may erode motivation and attention, weakening oversight effectiveness [1].
@@ 2.2 Attack Vectors and Hardening Strategies in Cybersecurity
With the implementation of AI systems across diverse fields of application, ensuring the security of deployed solutions has become increasingly urgent. Achieving this requires an in-depth understanding of specific attack vectors and the attack surface within each application domain [30]. Traditionally, an attack vector is defined as a method or pathway that a malicious actor uses to gain access to a network or system. In this paper, we extend the definition to also include methods or pathways that undermine the effectiveness of human oversight. We use the term attack surface to describe the total set of points where a malicious actor could attempt to enter, exploit, or otherwise attack digital and AI infrastructures [30]. For example, research has systematically analyzed attack vectors in autonomous driving systems and defined the corresponding attack surface [31]. Other work has examined the attack surface in federated learning workflows [32]. For human oversight, however, such systematic investigations of the attack surface are still missing.
@@
Relevant related work can be found in the field of AI cybersecurity, which examines attack vectors that explicitly target AI systems. For instance, poisoning attacks manipulate training data to introduce backdoors into AI models that can later be exploited [33, 34, 35]. Adversarial attacks embed imperceptible patterns into inputs to change model behavior, often causing misclassifications or other problematic outcomes [36, 37, 38]. Another line of research targets explainable AI (XAI) methods designed to enhance understanding of AI decisions [25]. Such attacks manipulate explanations to obscure the real reasons behind a model’s decision [39, 40, 41].
Because human oversight extends beyond an isolated AI system, additional relevant research comes from traditional cybersecurity. Distributed denial-of-service (DDoS) attacks flood networks with traffic, making them unresponsive or unavailable [42, 43]. Man-in-the-middle attacks intercept and manipulate data traffic between two endpoints [44]. Another common vector is malware, which – due to its diversity – can be deployed against nearly any digital system [45, 46].
Particularly relevant for defining the attack surface are attacks on the human factor in cybersecurity. Social engineering exploits human vulnerabilities to extract sensitive information or gain access privileges [47, 48]. Two additional but in cybersecurity research less discussed vectors are coercion and bribery, which can likewise be used to gain sensitive information or access privileges [49, 50]. Finally, insider threats exploit individuals inside an organization (e.g., disgruntled employees) to carry out attacks [51]. With the growing threat of cyberattacks, every system that incorporates digital components must be analyzed for potential attack vectors. Our work contributes to this literature by contextualizing attack vectors for the specifics of human oversight of AI.
@@
In response to the many ways digital systems can be attacked, cybersecurity research has developed hardening strategies as methods designed to secure systems against attacks [52, 53, 54]. One line of work establishes methods for securing data traffic in computer networks. Network management systems can detect malicious traffic [55, 56], while encryption protects legitimate communications [57, 58]. To monitor organizational IT systems, intrusion detection systems have been introduced and are continuously improved [59, 60]. Another widely used practice is red teaming, where every system component is rigorously tested by simulating attacks; research continues to optimize red teaming approaches [61, 62]. Beyond technical strategies, organizational measures also play a key role. Transparency of software components is crucial for analyzing systems and identifying potentially harmful functionalities [63, 64]. Other well-established approaches include governance principles and audit frameworks that help secure deployed systems [65, 66]. Finally, like attack vectors, hardening strategies must also address the human factor. Research has explored how to improve system security by training users and raising awareness of risks [67, 68].
@@
In this paper, we extend the literature by contextualizing hardening strategies specifically for securing human oversight of AI against malicious attacks.
3 The Human Oversight Attack Surface
@@
Human oversight can serve as part of the safety, security, and accountability architecture of AI systems, as oversight personnel are expected to detect and mitigate risks. In addition, oversight personnel may be required to actively confirm that an AI system is in a safe state. Human oversight may become even more important in AI operations if it evolves into a central requirement of AI governance and regulation in high-risk contexts [2, 6, 3, 1].

### v2

# Secure human oversight of AI: Threat modeling in a socio-technical context

@@ Abstract
Human oversight of AI is promoted as a safeguard against risks such as inaccurate outputs, system malfunctions, or violations of fundamental rights, and is mandated in regulation like the European AI Act. Yet debates on human oversight have largely focused on its effectiveness, while overlooking a critical dimension: the security of human oversight. We argue that human oversight creates a new attack surface within the safety, security, and accountability architecture of AI operations. Drawing on cybersecurity perspectives, we model human oversight as an IT application for the purpose of systematic threat modeling of the human oversight process. Threat modeling allows us to identify security risks within human oversight and points towards possible mitigation strategies. Our contributions are: (1) introducing a security perspective on human oversight, (2) offering researchers and practitioners guidance on how to approach their human oversight applications from a security point of view, and (3) providing a systematic overview of attack vectors and hardening strategies to enable secure human oversight of AI.
@@ 1 Introduction
Human oversight of artificial intelligence (AI) is often considered a strategy to address risks associated with the use of AI in high-risk contexts [25, 45, 67]. Human oversight personnel are expected to detect inaccurate outputs, outputs that may pose risks to fundamental human rights (e.g., discriminatory outputs), unsafe states of AI systems, and system malfunctions, and to intervene in order to mitigate or minimize these risks [41]. Human oversight is also a key concept in AI governance and ethics [45, 25] and a central requirement for the implementation of high-risk AI systems in emerging AI regulation, such as Article 14 of the European AI Act [18, 45]. However, relying on humans to oversee AI also faces criticism and substantial challenges in practice [34, 25, 72]. The most prominent criticisms and challenges include doubts about whether humans can effectively oversee AI systems [25], the difficulty of defining what constitutes effective human oversight [72], clarifying who should provide oversight and when it should occur [18], specifying which risks and unwanted system behaviors oversight personnel should detect [41], and determining how to design and evaluate effective human oversight [72].
@@
However, one key challenge that remains underexplored is the security risks that arise from human oversight of AI. Because human oversight will be implemented as part of the safety, security, and accountability architecture of AI systems [45, 34], malicious actors who successfully attack elements related to oversight can significantly undermine the safety of AI operations. In this sense, the human oversight architecture creates a new attack surface for targeting AI operations in high-risk contexts such as critical infrastructure, public administration, or medicine. A security-focused perspective on human oversight is crucial, given the central role human oversight already plays, the expectations placed on it in high-risk contexts, and its growing importance as a regulatory requirement in AI governance, such as under the EU AI Act. If we fail to assess and address the security challenges of human oversight, we risk overlooking potential attack vectors in high-risk contexts; and without awareness of these vectors, we will not be able to design effective countermeasures and hardening strategies. Ultimately, neglecting security aspects of human oversight can undermine its main objective—risk mitigation—or even increase risks by introducing additional attack vectors.
In this paper, we introduce a security-related perspective on human oversight of AI. Human oversight is best understood as a socio-technical system that can be instantiated in many different ways. For the purposes of threat modeling, we model it as an IT application and draw on established cybersecurity methods to analyze attack surfaces and vulnerabilities. This enables us to identify security threats to the key requirements of effective human oversight (i.e., epistemic access, self-control, causal power, and fitting intentions; [72]). Threat modeling further provides insights into applicable countermeasures and hardening strategies to mitigate the identified cybersecurity risks.
@@
This paper makes three key contributions. First, we introduce a security-related perspective on human oversight of AI. This perspective is critically missing from research, even though human oversight is becoming an integral component of the safety, security, and accountability architecture of AI operations, and thus an increasingly viable target for malicious actors. Second, we provide guidance on how to perform threat modeling in socio-technical contexts and identify cybersecurity threats of human oversight with possible countermeasures. Thereby, this work provides a starting point for future threat modeling of implementations of human oversight. Third, we provide a list of security vulnerabilities and hardening strategies that are applicable to human oversight applications in general and should always be considered by practitioners. We hope these contributions support AI governance and system design in ensuring not only effective but also secure human oversight of AI.
@@ 2.1 Effective Human Oversight of AI
Research discusses the nature of human oversight of AI, including the what, when, and who of oversight. Regarding the what, many forms of human-AI collaboration involve human oversight. Across a wide range of human-AI interaction use cases, humans monitor AI system outputs, check their accuracy, request explanations, adapt outputs, intervene when systems malfunction, or take over manual control in the event of failures [40, 10, 81, 17, 47, 73, 46, 13]. In a more targeted sense, and following how emerging research has defined the concept [72], human oversight specifically refers to humans mitigating risks in AI operations [50, 18]. More concretely, oversight personnel should monitor AI systems, detect failures and malfunctions, identify unsafe states, flag inaccurate outputs, and recognize outputs that may pose risks to fundamental human rights, such as discriminatory outputs [72, 41, 18]. Once such issues are detected, oversight personnel are expected to act in ways that effectively mitigate risks [41, 72]. This can involve taking over manual control, overwriting inaccurate or inadequate outputs [14], or stopping AI operations and triggering intervention by other human stakeholders, such as developers who can adapt the system. Importantly, the AI system under oversight may be operated by another human who is not considered oversight personnel but a user of the system. For example, a patient using an AI chatbot for mental health support [76] would be the user, while oversight personnel monitor the patient–AI interaction to detect risky situations (e.g., if the patient discusses suicidal behavior with the chatbot).
@@
Regarding the when, human involvement is important at every stage of AI development: during system design, at run time, and when inspecting errors and failures in hindsight [67, 68, 24]. In a more targeted sense, human oversight usually refers to personnel overseeing AI systems during run time [72, 45]. This ranges from dynamic, time-critical situations like autonomous vehicle oversight [14, 39] to less time-pressured contexts like judicial AI [41, 33].
@@
Regarding the who of human oversight, any human with the skills necessary to effectively oversee an AI system could serve as oversight personnel [18, 78]. These individuals may be domain experts such as doctors or judges, who can evaluate the adequacy of specific AI outputs based on their expertise. Alternatively, oversight personnel may be trained to detect AI failures without being domain experts in the task performed by the AI. For example, in a medical AI setting, one oversight person might monitor the system’s technical functions on an aggregate level, while medical professionals evaluate individual outputs requiring domain knowledge [78, 41]. Human oversight can also be performed by a single individual or by a team collaboratively overseeing AI systems. Research has also examined the requirements for effective human oversight. Sterz et al. [72] identify four key aspects: epistemic access, causal power, self-control, and fitting intentions.
@@
Epistemic access means that oversight personnel possess sufficient knowledge about the AI system and understand the consequences of their interventions. Without it, they may fail to detect system failures or inaccurate outputs. Causal power refers to the ability to intervene in ways that influence AI processes and outputs. Without it, oversight personnel may be unable to overwrite or correct system outputs. Self-control requires oversight personnel to maintain agency over their own actions. For instance, fatigue or intoxication may impair their capacity to fulfill oversight duties. Likewise, being threatened by another person can also undermine self-control, as the actions performed by the oversight personnel are then steered by another agent. Fitting intentions means that personnel are motivated to carry out their responsibilities with the right intentions. In an extreme case, a malicious actor might be hired as oversight personnel and deliberately undermine AI safety. In a less extreme case, poor working conditions may erode motivation and attention, weakening oversight effectiveness [72].
@@ 2.2 Attack Vectors and Hardening Strategies in Cybersecurity
With the implementation of AI systems across diverse context, ensuring the security of deployment has become increasingly important. Enabling systematic threat modeling for AI-based applications requires an in-depth understanding of specific attack vectors and the attack surface within an application domain [74]. Traditionally, an attack vector is defined as a method or pathway that a malicious actor uses to gain access to a network or system. In this paper, we extend the definition to also include methods or pathways that undermine the effectiveness of human oversight. We use the term attack surface to describe the total set of points where a malicious actor could attempt to enter, exploit, or otherwise attack digital and AI infrastructures [74]. While there is research that systematically analyses AI-based systems from a cybersecurity point of view [22, 60], such systematic investigations are still missing for human oversight.
@@
Related work can be found in the field of AI cybersecurity, which examines attack vectors that explicitly target AI systems. For instance, poisoning attacks manipulate training data to introduce backdoors into AI models that can later be exploited [65, 26, 32]. Adversarial attacks embed imperceptible patterns into inputs to change model behavior, often causing misclassifications or other problematic outcomes [29, 83, 2]. Another line of research targets explainable AI (XAI) methods designed to enhance understanding of AI decisions [42]. Such attacks manipulate explanations to obscure the real reasons behind a model’s decision [70, 1, 69]. However, AI system are still IT systems at the core. This means that traditional attack vectors, like distributed denial-of-services (DDOS) [43, 58], man-in-the-middle (MITM) [15], or social engineering attacks [5, 16], have to be considered for AI systems as well. Furthermore, human-centered attack vectors that are usually underexplored in cybersecurity, e.g., coercion, bribery, and insider threats like disgruntled employees [28, 75, 48], are of special importance for human oversight.
@@
In response to the many ways digital systems can be attacked, cybersecurity research has developed hardening strategies as methods designed to secure AI systems against attacks [62, 61, 9]. One line of work establishes methods for securing data traffic in computer networks. Network management systems can detect malicious traffic [71, 31], while encryption protects legitimate communications [52, 3]. To monitor organizational IT systems, intrusion detection systems have been introduced and are continuously improved [36, 35]. Another widely used practice is red teaming, where every system component is rigorously tested by simulating attacks; research continues to optimize red teaming approaches [54, 82]. Beyond technical strategies, organizational measures also play a key role. Transparency of software components is crucial for analyzing systems and identifying potentially harmful functionalities [30, 27]. Other well-established approaches include governance principles and audit frameworks that help secure deployed systems [63, 19]. Finally, like attack vectors, hardening strategies must also address the human factor. Research has explored how to improve system security by training users and raising awareness of risks [23, 55].
@@
In this paper, we extend the literature by contextualizing attack vectors and hardening strategies specifically for securing human oversight of AI and show how to use this knowledge for systematic threat modeling of human oversight.
3 Threat modeling of human oversight
@@
Human oversight can serve as part of the safety, security, and accountability architecture of AI systems, as oversight personnel are expected to detect and mitigate risks. In addition, oversight personnel may be required to actively confirm that an AI system is in a safe state. Human oversight may become even more important in AI operations if it evolves into a requirement of AI governance and regulation in high-risk contexts [25, 18, 45, 72].